Differentiate the following with regard to x :

y=x^4sinx


1
Expert's answer
2021-07-06T15:39:52-0400

Given the equation


"y=x^{4sinx}"

The derivative of the function is obtained thus:


"\begin{aligned}
&\begin{gathered}
\quad \frac{\mathrm{d}}{\mathrm{d} x}\left[x^{4 \sin (x)}\right] \\
=x^{4 \sin (x)} \cdot \frac{\mathrm{d}}{\mathrm{d} x}[\ln (x) \cdot 4 \sin (x)] \\
=x^{4 \sin (x)} \cdot 4 \cdot \frac{\mathrm{d}}{\mathrm{d} x}[\ln (x) \sin (x)] \\
=4 x^{4 \sin (x)}\left(\frac{\mathrm{d}}{\mathrm{d} x}[\ln (x)] \cdot \sin (x)+\ln (x) \cdot \frac{\mathrm{d}}{\mathrm{d} x}[\sin (x)]\right) \\
=4 x^{4 \sin (x)}\left(\frac{1}{x} \sin (x)+\ln (x) \cos (x)\right) \\
=4 x^{4 \sin (x)}\left(\frac{\sin (x)}{x}+\cos (x) \ln (x)\right)
\end{gathered}\\
&\text { We can rewrite the above result as: }\\
&=x^{4 \sin (x)}\left(\frac{4 \sin (x)}{x}+4 \cos (x) \ln (x)\right)
\end{aligned}"